# Per-Gaussian Embedding-Based Deformation for Deformable 3D Gaussian Splatting

### v1

@@ 5 Conclusion and Limitation
We propose a per-Gaussian deformation for 3DGS that takes per-Gaussian embeddings as input instead of using the typical deformation field of deformable 3DGS methods, achieving high performance. We enhance the quality by decomposing the dynamic changes into coarse and fine deformation. However, our method also has limitations. The dynamic regions of our rendering results tend to be blurred when there is significant motion between frames, as in the other baselines (Figure 10). It is expected to be resolved with additional supervision like correspondences in rapidly changing dynamic regions. In addition, our method tends to be slower in rendering compared to existing Gaussian Splatting methods. This could be improved by skipping the deformation prediction for static areas and pruning unnecessary points after training, as mentioned before.
Acknowledgments
@@
Appendix 0.A More Analysis
0.A.1 Effect of multi-view DSSIM loss
As shown in Figure S1, we observe that DSSIM loss helps to learn difficult areas in the scene, such as the details of the background. However, optimizing DSSIM loss at every step [8] is computationally expensive, and the introduction of DSSIM loss increases the number of Gaussians, further extending the training time. Our multi-view DSSIM loss achieves a similar level of quality to what is computed at every step while reducing the number of computations for DSSIM.
0.A.2 Effect of embedding size
@@
| 256 / 16 | 32.79 | 0.960 | 0.030 | 128 / 32 | 33.23 | 0.962 | 0.028 |
| 256 / 32 | 33.28 | 0.963 | 0.028 | 256 / 32 | 33.28 | 0.963 | 0.028 |
| 256 / 64 | 32.60 | 0.960 | 0.031 | 512 / 32 | 32.80 | 0.961 | 0.031 |
We report the performance variation by changing the size of per-Gaussian and temporal embeddings in Table S1. We observe that the model is sensitive to the size of while it is less sensitive to the size of . We choose the dimensions of and to be 32 and 256, respectively.
Appendix 0.B Training Details
0.B.1 Details of Efficient Training Strategy
For camera sampling, we exclude cameras close to the previous camera center for the next sampling candidates. This is achieved by calculating the pairwise distance between all camera origin and excluding the cameras below the 40 percentile of distances.
For frame sampling, we record the L1 error between rendered images and ground truth for each frame during training. We treat this as a probability distribution for frame sampling to focuse on frames with higher errors. For the first 10K iterations, we randomly sample the frames, and then alternately use random sampling and error-based frame sampling.
Finally, we use multi-view DSSIM loss with L1 loss to induce densification in areas where the model struggles, such as background or dynamic regions. Due to the computational cost of SSIM, we fix the frame obtained by loss-based sampling for every 50 iterations, and minimize the multi-view DSSIM loss through camera sampling during the next 5 iterations. Similar to frame sampling, DSSIM loss is applied after the first 10K iterations.
0.B.2 Implementation Details
@@
In all experiments, we use a 32-dimensional vector for Gaussian embeddings and a 256-dimensional vector for temporal embeddings . The decoder MLP consists 128 hidden units and the MLP head for Gaussian parameters are both composed of 2 layers with 128 hidden units. For Technicolor and HyperNeRF datasets, we use a 1-layer decoder MLP. To efficiently and stabilize the initial training, we start the temporal embedding grid at the same time resolution as and gradually increase it to resolution over 10K iterations, with set to 150 for 300 frames. The learning rate of our deformation decoder starts at and exponentially decays to over the total training iterations. The learning rate of the temporal embedding follows that of the deformation decoder, while the learning rate of per-Gaussian embedding is set to . We eliminate the opacity reset step of 3DGS and instead add a loss that minimizes the mean of Gaussian opacities with a weight of . We empirically set the start of the training strategy to 10K iterations. Following Sync-NeRF, we introduce a learnable time offset to each camera in Neural 3D Video and Technicolor Light Field dataset.
@@
For initialization points, we downsample point clouds obtained from COLMAP dense scene reconstruction. For Neural 3D Video dataset, we train for 80K iterations; for Technicolor dataset, 100K for the Birthday and Painter scenes and 120K for the rest; for the HyperNeRF dataset, 80K for the banana scene and 60K for the others. We periodically use the DSSIM loss at specific steps, excluding the HyperNeRF dataset. For calculating computational cost, the Neural 3D Video dataset was measured using A6000, while the Technicolor Light Field Dataset and the HyperNeRF Dataset were measured using 3090.
Appendix 0.C More Results
@@
Table S2-S4 report the quantitative results of each scene on Neural 3D Video Dataset, Technicolor Light Field Dataset, and HyperNeRF Dataset. In most scenes, our model shows better perceptual performance than the baselines. Additionally, we show the entire rendering results in Figure S2-S4.
0.C.1 Per-Scene Results of Neural 3D Video Dataset
@@
| 4DGS | 28.21 | 0.913 | 28.68 | 0.917 | 28.45 | 0.913 | 33.55 | 0.961 | 34.02 | 0.963 |
| 4DGaussians | 29.28 | 0.929 | 29.40 | 0.927 | 29.13 | 0.925 | 33.90 | 0.961 | 33.37 | 0.960 |
| Ours | 29.76 | 0.931 | 30.12 | 0.935 | 30.28 | 0.934 | 33.55 | 0.963 | 33.55 | 0.963 |
0.C.2 Per-Scene Results of Technicolor Light Field Dataset
@@
| Ours | 36.63 | 0.923 | 30.64 | 0.864 | 31.84 | 0.922 |
0.C.3 Per-Scene Results of HyperNeRF Dataset

### v2

@@ 5 Conclusion and Limitation
We propose a per-Gaussian deformation for 3DGS that takes per-Gaussian embeddings as input, instead of using the typical deformation fields from previous deformable 3DGS works, resulting in high performance. We enhance the reconstruction quality by decomposing the dynamic changes into coarse and fine deformation. However, our method learns inappropriate Gaussian deformation with casually captured monocular videos[7], like other baselines. We plan to address it in future work by introducing useful prior for monocular video settings.
Acknowledgements
@@
Appendix 0.A Training Details
0.A.1 Efficient Training Strategy
We introduce a training strategy for faster convergence and higher performance. The first is to evenly cover multi-view camera perspectives and exclude the camera index that was used in the previous iteration. We pre-compute pairwise distances of all camera origins before the start of the training. Then we exclude the cameras with distance to the previous camera less than 40 percentile of all distances.
The second is to sample frames within the target cameras. Different frames may have different difficulties for reconstruction. We measure the difficulty of a frame by the magnitude of error at the frame. Then, we sample training frames with a categorical distribution which is proportional to the magnitude of error. For multi-view videos, we average the errors from all viewpoints. For the first 10K iterations, we randomly sample the frames, and then alternately use random sampling and error-based frame sampling.
The third is to change the policy for Gaussian densification. Previous methods invoke the densification by minimizing L1 loss [29] or L1 loss and DSSIM loss for every iteration [31, 32, 13]. We observe that the DSSIM loss improves the visual quality in the background but takes longer training time. Therefore, we use L1 loss for every iteration and periodically use the additional multi-view DSSIM loss in the frame with high error. This encourages densification in the region where the model struggles. We fix the frame obtained by loss-based sampling for every 50 iterations, and minimize the multi-view DSSIM loss through camera sampling during the next 5 iterations. Similar to frame sampling, DSSIM loss is applied after the first 10K iterations.
0.A.2 Implementation Details
@@
| 256 / 16 | 33.53 | 0.962 | 0.030 | 128 / 32 | 33.61 | 0.964 | 0.029 |
| 256 / 32 | 33.57 | 0.964 | 0.028 | 256 / 32 | 33.57 | 0.964 | 0.028 |
| 256 / 64 | 33.71 | 0.964 | 0.029 | 512 / 32 | 33.55 | 0.964 | 0.028 |
We report the performance variation by changing the size of per-Gaussian and temporal embeddings in Table S1. The embedding size does not significantly affect the overall performance. In all experiments, considering the capacity and details, we use a 32-dimensional vector for Gaussian embeddings and a 256-dimensional vector for temporal embeddings .
@@
The decoder MLP consists 128 hidden units and the MLP head for Gaussian parameters are both composed of 2 layers with 128 hidden units. For Technicolor and HyperNeRF datasets, we use a 1-layer decoder MLP. To efficiently and stabilize the initial training, we start the temporal embedding grid at the same time resolution as and gradually increase it to resolution over 10K iterations, with set to 150 for 300 frames. The learning rate of our deformation decoder starts at and exponentially decays to over the total training iterations. The learning rate of the temporal embedding follows that of the deformation decoder, while the learning rate of per-Gaussian embedding is set to . We eliminate the opacity reset step of 3DGS and instead add a loss that minimizes the mean of Gaussian opacities with a weight of . We empirically set the start of the training strategy to 10K iterations.
For the Neural 3D Video dataset, we follow Sync-NeRF[11] to introduce trainable time offsets to each camera. In addition, we learn the deformation of color, scale, and opacity from 5K iterations and perform densification and pruning.
@@
For initialization points, we downsample point clouds obtained from COLMAP dense scene reconstruction. For the Neural 3D Video dataset, we train for 80K iterations; for the Technicolor dataset, 100K for the Birthday and Painter scenes and 120K for the rest; for the HyperNeRF dataset, 80K for the Banana scene and 60K for the others. We periodically use the DSSIM loss at specific steps, excluding the HyperNeRF dataset. For calculating computational cost, the Neural 3D Video dataset was measured using A6000, while the Technicolor Light Field dataset and the HyperNeRF dataset were measured using 3090.
For local smoothness embedding regularization, we set the weight of loss to 1 on the Neural 3D Video dataset and 0.1 on the Technicolor Light Field dataset and the HyperNeRF dataset.
Appendix 0.B More Results
@@
We show the additional rendering of 4DGasussians [29] in Figure S2. And, we show the entire rendering results of ours in Figure S3-S5. In most scenes, our model shows better perceptual performance than the baselines. Table S2-S4 report the quantitative results of each scene on the Neural 3D Video dataset, Technicolor Light Field dataset, and HyperNeRF dataset.
0.B.1 Additional Results of 4DGaussians
The results of 4DGaussians are produced using the official repository. In the teaser, we choose the third fragment of flame salmon scene which contains more difficult movements than the typical first fragment to demonstrate our superiority.
0.B.2 Per-Scene Results of Neural 3D Video Dataset
@@
| 4DGaussians | 28.21 | 0.913 | 28.68 | 0.917 | 28.45 | 0.913 | 33.55 | 0.961 | 34.02 | 0.963 |
| 4DGS | 29.28 | 0.929 | 29.40 | 0.927 | 29.13 | 0.925 | 33.90 | 0.961 | 33.37 | 0.960 |
| Ours | 29.69 | 0.934 | 29.94 | 0.934 | 29.90 | 0.933 | 33.57 | 0.964 | 33.45 | 0.963 |
0.B.3 Per-Scene Results of Technicolor Light Field Dataset
@@
| Ours | 36.18 | 0.924 | 31.07 | 0.868 | 31.33 | 0.912 |
0.B.4 Per-Scene Results of HyperNeRF Dataset